# A Bayesian Dirichlet Auto-Regressive Conditional Heteroskedasticity Model for Compositional Time Series

### v1

@@ 3.2.3 Simulation Studies 3 & 6: tVARMA DGP
$\boldsymbol{\eta}_{t}\sim\mathcal{N}(\boldsymbol{\beta}+\boldsymbol{A}(\text{alr}(\mathbf{y}_{t-1})-\boldsymbol{\beta}),\sigma^{2}\boldsymbol{\Sigma})$
$\mathbf{y}_{t}=\text{alr}^{-1}(\boldsymbol{\eta}_{t}).$
A single precision parameter $\sigma$ was drawn independently for each simulation from $\sigma\sim\mathcal{U}(0.05,0.5).$ The covariance matrix $\boldsymbol{\Sigma}$ was generated to be positive-definite with complex interdependencies $\boldsymbol{\Sigma}=\mathbf{M}^{\top}\mathbf{M},$ where $\mathbf{M}$ was a $4\times 4$ matrix with elements drawn independently from $\mathcal{U}(-0.3,0.3)$ for each simulation.
3.2.4 Simulated Shocks
In simulations 1-3, random shocks were introduced into the training data to simulate misreported observations. Time points were selected by sampling intervals from a Poisson(6) distribution and cumulatively summing them. At these times $l$, $\mathbf{y}_{t}$ was replaced with a random composition generated by independently drawing each component from a Uniform(0,1) distribution and normalizing the vector to sum to 1,
$y_{lj}\sim\mathcal{U}(0,1),\ \mathbf{y}_{\text{misreported}[l]}=\frac{y_{lj}}{\sum_{j=1}^{5}y_{lj}},j=1,\dots,5.$
3.2.5 Regime Shifts
In Simulation 4-6, we introduced regime shifts at a random time point $t_{\text{shift}}$ uniformly selected from [10, 50] by redrawing the intercept vector $\boldsymbol{\beta}$, the auto-regressive matrix $A$ and the precision parameters (either $\phi_{0}$, $\sigma$, $\Sigma$) for the DARMA/DARCH models or tVARMA model. For $t=t_{\text{shift}}$ to $t=t_{\text{shift}}+9$, data were generated using these new parameters, simulating a regime shift in the data-generating process. After this period, parameters reverted to their original values, allowing us to evaluate the models’ robustness to temporary structural changes.
3.3 Model Comparisons
We fitted three models to each simulated dataset for comparison: the B-DARMA(1,0)-DARCH(1,1), the Bayesian transformed VARMA (B-tVARMA(1,0)), and the B-DARMA(1,0) models. All models shared the following priors: the intercepts $\boldsymbol{\beta}$ were given $\text{Normal}(0,0.3^{2})$ priors, and the elements of the auto-regressive matrices $\boldsymbol{A}_{p}$ had $\text{Normal}(0,1)$ priors.
For the B-DARMA(1,0)-DARCH(1,1) model, the initial precision parameter $\phi_{0}$ was assigned a $\text{Normal}(7,1.5^{2})$ prior, and the AR and MA terms for $\phi_{t}$ had $\text{Normal}(0.35,0.5^{2})$ and $\text{Normal}(-0.75,0.5^{2})$ priors, respectively.
@@ 3.3 Model Comparisons
All models were fit using Stan with four chains of 1,000 iterations each (500 warm-up), yielding 2,000 posterior samples.
@@ 3.4 Metrics
Additionally, we assess the Partial Autocorrelation Function (PACF) (see Supplementary Section Definition of the Partial Autocorrelation Function (PACF)) of the sums of squared standardized residuals to evaluate how well each model captures residual dynamics during the in-sample period. For each model, residuals are standardized according to the model’s underlying distribution (see Supplementary Section Standardized Residuals).
@@ 3.5 Results
3.5.1 Simulation Studies 1-3: Shocks
The left side of table 1 shows the forecast performance FRMSE and FMAE metrics for the B-DARMA, B-DARCH, and B-tVARMA models across 50 simulated datasets in three different scenarios.
Across simulation studies 1–3, the B-DARCH model consistently had the lowest FRMSE and FMAE values compared to the second-best model, the B-tVARMA. In simulation study 1, FRMSE and FMAE were 1.6% and 1.4% lower, respectively. In simulation study 2, they were 12.6% and 7.7% lower. In simulation study 3, FRMSE and FMAE were .8% and 3.0% lower than B-tVARMA.
Figure S1 shows the mean PACF values across all simulations for B-DARCH, B-DARMA, and B-tVARMA models. The mean PACF values for the B-DARCH model consistently remain closer to zero across the majority of lags, with fewer pronounced PACF values. Notably, the B-DARMA and B-tVARMA models show large negative PACF values, particularly for early lags (lags 1-5). These large values indicate that residuals in these models have substantial autocorrelation, suggesting less efficient capture of the underlying dependencies across time within the data. In contrast, B-DARCH shows minimal autocorrelation for early lags, reflecting an overall improvement in residual behavior and a better fit to the true data-generating process.
3.5.2 Simulation Studies 4-6: Regime Shifts
The right side of Table 1 shows the forecast performance metrics FRMSE and FMAE for the B-DARMA, B-DARCH, and B-tVARMA models across 50 simulations with regime shifts. The B-DARCH model consistently had the lowest FRMSE and FMAE values, outperforming the second-best model (B-tVARMA) with 14.6% lower FRMSE and 8.4% lower FMAE in simulation study 4, 4.7% and 4.6% lower in simulation study 5, and 10.5% and 4.8% lower in simulation study 6. These results highlight the B-DARCH model’s comparative advantage in adapting to sudden changes in the underlying dynamics.
Figure S2 shows the mean PACF values for simulations 4-6 for B-DARCH, B-DARMA, and B-tVARMA models. The B-DARCH model has positive peaks for early lags, with autocorrelation generally settling closer to zero across subsequent lags. Notably, the magnitude of the PACF values for B-DARCH is smaller than those of the other models, except for Simulation 4, where the B-DARMA model shows a smaller magnitude at lag 1.
In contrast, both B-DARMA and B-tVARMA have strong positive PACF values for the first few lags, which then gradually decrease to negative values at later lags, indicating greater persistence of autocorrelation effects. This pattern suggests that B-DARMA and B-tVARMA models struggle more with handling shifts in the underlying process, resulting in larger autocorrelations in early lags compared to the B-DARCH model.
These results indicate that the B-DARCH model’s dynamic volatility component enables it to adapt more effectively to structural breaks, capturing both systematic and stochastic changes in volatility. In contrast, the B-DARMA and B-tVARMA models, lacking a stochastic volatility structure, were less effective in handling the regime shifts, leading to higher forecast errors.
Model Simulation 1 (DARMA Shock) Simulation 4 (DARMA Regime Shift)
@@ 3.5.2 Simulation Studies 4-6: Regime Shifts
B-DARMA 38.6 2.94 25.3 2.58
B-DARCH 38.0 2.87 18.1 2.19
B-tVARMA 38.6 2.91 21.2 2.37
Model Simulation 2 (DARCH Shock) Simulation 5 (DARCH Regime Shift)
@@
B-DARMA 11.2 1.84 20.4 2.28
B-DARCH 9.79 1.69 18.3 2.09
B-tVARMA 11.3 1.83 19.2 2.19
Model Simulation 3 (tVARMA Shock) Simulation 6 (tVARMA Regime Shift)
Average FRMSE Average FMAE Average FRMSE Average FMAE
B-DARMA 25.6 2.43 32.5 2.59
B-DARCH 24.1 2.24 22.9 2.18
B-tVARMA 24.3 2.31 25.7 2.29
Table 1: Summary of Model Performance Metrics times 100 on test set ($T=40$) across 50 simulations for Simulation Studies 1–6. Average FRMSE is the average Forecasted Root Mean Squared Error across all five components; Average Forecasted Mean Absolute Error is the average FMAE across all five components.
@@ 4.2 Objective
The primary goal of this analysis is to compare the performance of three model classes: Bayesian Dirichlet Auto-Regressive Moving Average (B-DARMA), Bayesian transformed Vector Auto-Regressive Moving Average (B-tVARMA), and Bayesian Dirichlet Auto-Regressive Moving Average with Dirichlet Auto-Regressive Conditional Heteroskedasticity (B-DARMA-DARCH).
@@ 4.3 Comparison Models
The B-DARCH and B-DARMA models share the same model for $\boldsymbol{\eta}$ but differ in their treatment of the precision parameter. In the B-DARCH model, the precision parameter has its own AR and MA components, allowing for dynamic heteroskedasticity that evolves over time. In contrast, the B-DARMA model does not include these components in the precision parameter, relying solely on auto-regressive and moving average terms in the mean structure to account for time dependence.
The third model, B-tVARMA, is a Bayesian normal VARMA model applied to transformed data. It transforms the observed multivariate time series $\mathbf{y}_{t}$ using the additive log-ratio (alr) transformation, resulting in ${\textnormal{alr}}(\mathbf{y}_{t})\in\mathbb{R}^{J-1}$. The transformed data is then modeled as ${\textnormal{alr}}(\mathbf{y}_{t})\sim N(\boldsymbol{\eta}_{t},\mathbf{\Sigma})$, where $\boldsymbol{\eta}_{t}$ has time-varying mean defined in (4), and $\bf{\Sigma}$ is an unknown $J-1\times J-1$ positive definite covariance matrix. The covariance matrix $\bf{\Sigma}$ is modeled through its Cholesky factor $L_{\Sigma}$, scaled by $\sigma$, $\bf{\Sigma}$ $=\sigma^{2}\Omega$.
All models are fit with STAN (Stan Development Team, 2022) in R. We run $4$ chains with $2000$ iterations each with a warm up of $1000$ iterations for a posterior sample of $4000$. Initial values for all parameters are selected randomly from the interval $[-1,1]$.

### v2

@@ 3.2 Simulation Setup
$\boldsymbol{\eta}_{t}\;=\;\mathbf{A}_{t,1}\!\big(\mathbf{Y}_{t-1}-\mathbf{X}_{t-1}\boldsymbol{\beta}\big)\;+\;\mathbf{X}_{t}\boldsymbol{\beta},\qquad\mathbf{Y}_{t}\mid\boldsymbol{\eta}_{t}\sim\mathcal{N}(\boldsymbol{\eta}_{t},\,\boldsymbol{\Sigma}),\quad\boldsymbol{\Sigma}=\operatorname{diag}(\boldsymbol{\sigma})\,\boldsymbol{\Omega}\,\operatorname{diag}(\boldsymbol{\sigma}).$
The coefficient matrix follows a mean‑reverting AR(1) evolution around a long‑run target:
$\mathbf{A}_{t,1}\;=\;\bar{\mathbf{A}}_{1}\;+\;\rho^{(A)}_{1}\!\big(\mathbf{A}_{t-1,1}-\bar{\mathbf{A}}_{1}\big)\;+\;\tau^{(A)}\,\boldsymbol{Z}^{(A)}_{t}$,
with $\operatorname{vec}(\boldsymbol{Z}^{(A)}_{t})\stackrel{{\scriptstyle\text{iid}}}{{\sim}}\mathcal{N}(\mathbf{0},\mathbf{I})$ and $\tau^{(A)}\sim\mathrm{Half\!-\!Normal}(0,\,0.1^{2})$.
3.3 Data-Generating Processes (DGPs)
Studies 1 & 3: DARMA
Compositions are drawn from a Dirichlet ARMA model with constant precision:
$\mathbf{Y}_{t}\mid\boldsymbol{\mu}_{t},\phi_{0}\sim\mathrm{Dirichlet}(\phi_{0},\boldsymbol{\mu}_{t}),\qquad\boldsymbol{\mu}_{t}\;=\;\mathrm{alr}^{-1}(\boldsymbol{\eta}_{t}),$
with $\phi_{0}\sim\mathcal{U}(6,\,7.5)$ independently per simulation.
Studies 2 & 4: tVARMA
On the ALR scale,
$\boldsymbol{\eta}_{t}\;\sim\;\mathcal{N}\!\big(\boldsymbol{\beta}+\mathbf{A}(\mathrm{alr}(\mathbf{Y}_{t-1})-\boldsymbol{\beta}),\ \boldsymbol{\Sigma}\big),\qquad\mathbf{Y}_{t}\;=\;\mathrm{alr}^{-1}(\boldsymbol{\eta}_{t}).$
For each simulation we set $\boldsymbol{\Sigma}=\mathbf{M}\mathbf{M}^{\top}$ with $M_{ij}\sim\mathcal{U}(-0.3,\,0.3)$ i.i.d., yielding diverse cross-component dependence.
3.4 Perturbations
Misreported observations.
During the training window, times are selected via cumulative sums of $\mathrm{Poisson}(6)$ draws. At such times $\ell$, replace the composition with a random simplex draw:
$y_{\ell j}\sim\mathcal{U}(0,1)\ \text{i.i.d.},\qquad\mathbf{Y}_{\text{misreported}[\ell]}\;=\;\frac{(y_{\ell 1},\dots,y_{\ell 5})}{\sum_{j=1}^{5}y_{\ell j}}.$
Temporary regime shifts.
@@ 3.4 Perturbations
3.5 Fitted Models and Priors
Each simulated dataset is fit with four Bayesian models: B-DARMA(1,0)-DARCH(1,1) (B-DARCH(1,1)) , B-DARMA(1,0), B-tVARMA(1,0), and B-TVP-VARMA(1,0). All models are implemented in Stan.
Priors.
Intercepts $\boldsymbol{\beta}\sim\mathcal{N}(0,\,0.3^{2})$. Elements of $\mathbf{A}_{p}$ in the fixed‑coefficient models receive elementwise $\mathcal{N}(0,\,1)$ priors. Gaussian‑error models use
$\boldsymbol{\Sigma}\;=\;\operatorname{diag}(\boldsymbol{\sigma})\,\boldsymbol{\Omega}\,\operatorname{diag}(\boldsymbol{\sigma}),\qquad L_{\Omega}\sim\mathrm{LKJ\_Corr\_Cholesky}(3),\quad\sigma_{c}\sim\mathrm{Half\!-\!Normal}(0,\,0.5^{2}).$
For B‑DARCH, the initial log‑precision $\phi_{0}\sim\mathcal{N}(7,\,1.5^{2})$ and the AR/MA terms in $\log\phi_{t}$ have $\mathcal{N}(0.35,\,0.5^{2})$ and $\mathcal{N}(-0.75,\,0.5^{2})$ priors, respectively.
For the B‑TVP–VARMA(1,0) comparator we impose mean reversion on the time‑varying coefficient matrix around a long‑run target. Specifically,
$\mathbf{A}_{t,1}\;=\;\bar{\mathbf{A}}_{1}\;+\;\rho^{(A)}_{1}\!\big(\mathbf{A}_{t-1,1}-\bar{\mathbf{A}}_{1}\big)\;+\;\tau^{(A)}\,\boldsymbol{Z}^{(A)}_{t},\quad\operatorname{vec}\!\big(\boldsymbol{Z}^{(A)}_{t}\big)\;\stackrel{{\scriptstyle\text{iid}}}{{\sim}}\;\mathcal{N}(\mathbf{0},\mathbf{I}),$
with priors
$\text{diag}(\bar{\mathbf{A}}_{1})\sim\mathcal{N}(0.5,\,0.3^{2}),\quad\bar{\mathbf{A}}_{1}^{(i\neq j)}\sim\mathcal{N}(0,\,0.2^{2}),\quad\rho^{(A)}_{1}\sim\operatorname{Beta}(9,1),\quad\tau^{(A)}\sim\mathrm{Half\!-\!Normal}(0,\,0.1^{2}),$
and shrinkage of the initial state to the target,
$\operatorname{vec}\!\big(\mathbf{A}_{1,1}-\bar{\mathbf{A}}_{1}\big)\sim\mathcal{N}(\mathbf{0},\,0.2^{2}\mathbf{I}).$
Computation.
We run four chains with 2,000 iterations per chain (1,000 warm‑up), yielding 4,000 post‑warm‑up draws per model. Convergence diagnostics and runtimes are reported in the supplement.
@@ 3.6 Metrics
Additionally, we assess the Partial Autocorrelation Function (PACF) (see Supplementary Section 6) of the sums of squared standardized residuals to evaluate how well each model captures residual dynamics during the in-sample period. For each model, residuals are standardized according to the model’s underlying distribution (see Supplementary Section 6).
@@ 3.7 Results
3.7.1 Simulation Studies 1–2: Shocks
Table 1 summarizes out-of-sample performance under (i) a DARMA DGP with misreported observations (Sim 1) and (ii) a tVARMA DGP with misreported observations (Sim 2). In both designs, B-DARCH attains the lowest average FRMSE and FMAE among all competitors. Relative to the second-best model, the FRMSE improvement is about $1.3\%$ in Sim 1 and $1.1\%$ in Sim 2; the corresponding FMAE improvements are about $2.7\%$ and $2.4\%$. The B-TVP-tVARMA comparator tracks B-tVARMA closely but remains slightly behind B-DARCH on both metrics.
Figure S1 displays mean PACF values of the sums of squared standardized residuals across replications. Under shock contamination, all models exhibit modest negative partial autocorrelation at the first few lags, consistent with one-step “correction” after a misreport. B-DARCH’s bars lie closest to zero for lags 1–6 and decay rapidly thereafter, indicating the least residual short-run dependence among the four methods.
3.7.2 Simulation Studies 3–4: Regime Shifts
Table 2 reports results when the DGP undergoes a temporary regime shift: (iii) DARMA with a transient parameter change (Sim 3) and (iv) tVARMA with a transient parameter change (Sim 4). B-DARCH again has the smallest average FRMSE and FMAE in both settings. The gains over the second-best model are modest but consistent: about $1.0\%$ (FRMSE) and $2.3\%$ (FMAE) in Sim 3, and about $0.2\%$ (FRMSE) and $0.9\%$ (FMAE) in Sim 4. Under a DARMA regime shift (Sim 3) the B-TVP-tVARMA comparator is the clear runner-up, whereas under a tVARMA regime shift (Sim 4) B-tVARMA is the runner-up.
Figure S2 shows the corresponding mean PACF patterns. With regime shifts, all models display a strong lag-1 spike; for both DGPs the spike is largest for the B-TVP-tVAR class (B-DARCH and B-TVP-tVARMA). Crucially, B-DARCH decays most rapidly and stays comparatively close to zero beyond lag 3, while B-tVARMA and B-TVP-tVARMA exhibit more persistent negative partial autocorrelations at medium lags (roughly 8–16), indicating longer-lived residual structure once the shift occurs. Taken together, these diagnostics suggest that B-DARCH concentrates any remaining dependence immediately after the shift and then dissipates it faster, whereas fixed-variance comparators leave more medium-horizon autocorrelation.
Summary.
Across all four designs (DARMA/tVARMA $\times$ shocks/regime shifts), B-DARCH yields the best point-forecast accuracy (FRMSE/FMAE) and the least residual persistence beyond very short lags. Because the DGPs considered here have constant variance, the improvements are necessarily modest; nevertheless, the volatility-aware specification helps the model absorb shocks and adapt quickly to transient parameter changes. Meanwhile, the B-TVP-tVARMA comparator is competitive, especially under regime shifts, but does not close the gap to B-DARCH.
Model Sim 1 (DARMA Shock) Sim 2 (tVARMA Shock)
@@ 3.7.2 Simulation Studies 3–4: Regime Shifts
B-DARMA 20.12 3.13 18.36 2.61
B-DARCH 19.67 2.98 18.08 2.53
B-tVARMA 19.97 3.07 18.28 2.59
B-TVP-tVARMA 19.92 3.06 18.33 2.61
Table 1: Summary of model performance metrics ($\times 100$) on the test set (T=40T{=}40) across 50 simulations for Simulation Studies 1–2. FRMSE ($\times 100$) equals the across-simulation average of $\mathrm{RMSE}=\sqrt{\frac{1}{TJ}\sum_{t=1}^{T}\sum_{j=1}^{J}(y_{tj}-\hat{y}_{tj})^{2}}$; FMAE ($\times 100$) equals the across-simulation average of $\mathrm{MAE}=\frac{1}{TJ}\sum_{t=1}^{T}\sum_{j=1}^{J}|y_{tj}-\hat{y}_{tj}|$.
Model Sim 3 (DARMA Regime Shift) Sim 4 (tVARMA Regime Shift)
@@
B-DARMA 18.68 2.72 18.11 2.48
B-DARCH 18.30 2.59 17.46 2.29
B-tVARMA 18.65 2.73 17.49 2.31
B-TVP-VAR 18.49 2.65 18.03 2.45
Table 2: Summary of model performance metrics ($\times 100$) on the test set (T=40T{=}40) across 50 simulations for Simulation Studies 3–4. FRMSE ($\times 100$) equals the across-simulation average of $\mathrm{RMSE}=\sqrt{\frac{1}{TJ}\sum_{t=1}^{T}\sum_{j=1}^{J}(y_{tj}-\hat{y}_{tj})^{2}}$; FMAE ($\times 100$) equals the across-simulation average of $\mathrm{MAE}=\frac{1}{TJ}\sum_{t=1}^{T}\sum_{j=1}^{J}|y_{tj}-\hat{y}_{tj}|$.
@@ 4.2 Objective
The primary goal is to compare four model classes for forecasting currency–share compositions: (i) Bayesian Dirichlet ARMA (B-DARMA), (ii) Bayesian Dirichlet ARMA with dynamic precision (B-DARMA-DARCH; “B-DARCH”), (iii) Bayesian transformed Gaussian VARMA (B-tVARMA), and (iv) a time-varying-parameter comparator in ALR space (B-TVP-tVARMA). All models are evaluated under a common training/validation/test split with weekly and yearly seasonality and a linear trend included in the design.
@@ 4.3 Comparison Models
B-DARCH (Dirichlet with DARCH precision).
B-DARCH and B-DARMA share the same ALR-scale mean model for $\boldsymbol{\eta}_{t}$, but B-DARCH endows the Dirichlet precision $\phi_{t}$ with its own AR/MA recursion on $\log\phi_{t}$, allowing time-varying dispersion (time‑varying volatility (conditional heteroskedasticity)) while keeping predictions on the simplex.
B-DARMA (Dirichlet with fixed precision regression).
Identical mean specification to B-DARCH, but the precision is deterministic given covariates $\mathbf{z}_{t}$ (no time‑varying volatility (conditional heteroskedasticity)). Temporal dependence is handled in the mean only.
B-tVARMA (transformed-Gaussian benchmark).
Transform the composition via $\mathbf{Y}_{t}={\textnormal{alr}}(\mathbf{y}_{t})\in\mathbb{R}^{J-1}$ and fit
$\mathbf{Y}_{t}\sim\mathcal{N}(\boldsymbol{\eta}_{t},\ \boldsymbol{\Sigma}),\qquad\boldsymbol{\eta}_{t}=\sum_{p=1}^{P}\mathbf{A}_{p}\!\big(\mathbf{Y}_{t-p}-\mathbf{X}_{t-p}\boldsymbol{\beta}\big)+\sum_{q=1}^{Q}\mathbf{B}_{q}\!\big(\mathbf{Y}_{t-q}-\boldsymbol{\eta}_{t-q}\big)+\mathbf{X}_{t}\boldsymbol{\beta},$
with $\boldsymbol{\Sigma}$ unknown. This model flexibly captures mean dynamics in ALR space but assumes time-invariant observation variance.
B-TVP-tVARMA (time-varying-parameter comparator in ALR space).
On the ALR scale we take
$\mathbf{Y}_{t}\mid\boldsymbol{\eta}_{t}\sim\mathcal{N}(\boldsymbol{\eta}_{t},\ \boldsymbol{\Sigma}),\qquad\boldsymbol{\Sigma}=\operatorname{diag}(\boldsymbol{\sigma})\,\boldsymbol{\Omega}\,\operatorname{diag}(\boldsymbol{\sigma}).$
The conditional mean includes both VAR and VMA components,
$\boldsymbol{\eta}_{t}=\sum_{p=1}^{P}\mathbf{A}_{t,p}\big(\mathbf{Y}_{t-p}-\mathbf{X}_{t-p}\boldsymbol{\beta}\big)+\sum_{q=1}^{Q}\mathbf{B}_{t,q}\big(\mathbf{Y}_{t-q}-\boldsymbol{\eta}_{t-q}\big)+\mathbf{X}_{t}\boldsymbol{\beta}.$
Coefficient matrices evolve by mean reversion toward long-run targets,
$\mathbf{A}_{t,p}=\bar{\mathbf{A}}_{p}+\rho^{(A)}_{p}\big(\mathbf{A}_{t-1,p}-\bar{\mathbf{A}}_{p}\big)+\tau^{(A)}_{p}\,\boldsymbol{Z}^{(A)}_{t,p},\quad\mathbf{B}_{t,q}=\bar{\mathbf{B}}_{q}+\rho^{(B)}_{q}\big(\mathbf{B}_{t-1,q}-\bar{\mathbf{B}}_{q}\big)+\tau^{(B)}_{q}\,\boldsymbol{Z}^{(B)}_{t,q}$,
with $\operatorname{vec}(\boldsymbol{Z}^{(A)}_{t,p})$ and $\operatorname{vec}(\boldsymbol{Z}^{(B)}_{t,q})$ i.i.d. $\mathcal{N}(\mathbf{0},\mathbf{I})$.
Computation.
All four models were fit in Stan with 4 chains (2,000 iterations; 1,000 warm‑up). Diagnostics were satisfactory across regions and models (no divergences; only rare max tree‑depth hits; worst $\hat{R}=1.01$; strong bulk/tail ESS); see Table S3.